# Towards foundational LiDAR world models with efficient latent flow matching

### v1

@@ Abstract
LiDAR-based world models offer more structured and geometry-aware representations than their image-based counterparts. However, existing LiDAR world models are narrowly trained; each model excels only in the domain for which it was built. Can we develop LiDAR world models that exhibit strong transferability across multiple domains? We conduct the first systematic domain transfer study across three demanding scenarios: (i) outdoor to indoor generalization, (ii) sparse-beam & dense-beam adaptation, and (iii) non-semantic to semantic transfer. Given different amounts of fine-tuning data, our experiments show that a single pre-trained model can achieve up to 11% absolute improvement (83% relative) over training from scratch and outperforms training from scratch in 30/36 of our comparisons. This transferability of dynamic learning significantly reduces the reliance on manually annotated data for semantic occupancy forecasting: our method exceed the previous semantic occupancy forecasting models with only 5% of the labeled training data required by prior models. We also observed inefficiencies of current LiDAR world models, mainly through their under-compression of LiDAR data and inefficient training objectives. To address this, we propose a latent conditional flow matching (CFM)-based frameworks that achieves state-of-the-art reconstruction accuracy using only half the training data and a compression ratio 6 times higher than that of prior methods. Our model achieves SOTA performance on future-trajectory-conditioned semantic occupancy forecasting while being 23x more computationally efficient (a 28x FPS speedup); and achieves SOTA performance on semantic occupancy forecasting while being 2x more computationally efficient (a 1.1x FPS speedup). Project Page: AdaFlowMatchingWM.github.io.
@@ 1 Introduction
World models enable agents to implicitly learn the dynamics of the environment by predicting future sensory observations, typically through generative models operating in a latent space rombach2022high ; ma2024latte . Recent advances have led to the development of numerous world models of RGB videos that have demonstrated impressive performance in applications such as autonomous driving russell2025gaia ; hu2023gaia ; wang2024driving , robotic navigation bar2024navigation , and other embodied tasks black2024pi_0 ; kim2024openvlaopensourcevisionlanguageactionmodel ; agarwal2025cosmos ; alhaija2025cosmos . These models can generate video sequences conditioned on historical frames and sometimes natural language input. While an image-based world model may suffice for repetitive tasks (e.g., learning robotic arm movements), its utility may be limited in complex tasks that demand geometrically structured information, such as autonomous driving. In these tasks, the lack of explicit semantic and geometric representations restricts practical applicability, as additional steps are still required to extract explicit spatial information.
@@
In contrast to RGB images, LiDAR provides rich geometric structure and implicitly gives a sparse representation for semantic cues about the environment. Unlike dense, pixel-based observations, a 3D object in a LiDAR point cloud is represented as a cluster of 3D points, encoding essential spatial and semantic information in a compact form. Despite these advantages, foundational world models of LiDAR data remain relatively underexplored: prior LiDAR world models are built for specific domains zhang2023copilot4d ; khurana2023point ; zhang2024bevworld . In comparison, recent advances in “foundation” models, e.g. pretraining with RGB image forecasting alhaija2025cosmos and trajectory forecasting zhou2024smartpretrain have demonstrated the effectiveness of pretraining in improving performance on downstream tasks. Can we develop a foundational LiDAR world model that leads to downstream performance gains on diverse forecasting tasks after fine-tuning? This goal is partly motivated by the fact that many of the causal factors that govern the motion of objects are shared across different domains and environments—many aspects of the dynamics of the world should, in principle, be transferable and can be observed in unlabeled LiDAR.
@@
Figure 1: Left: The overall pipeline of our method: we used the most readily publicly available LiDAR dataset from autonomous driving scenarios to train the proposed LiDAR world model. This well-trained world model is able to generalize well on the listed downstream tasks after fine-tuning, although scene and signal properties are quite different. Right: The comparison of our proposed world model with previous methods on nuScenes 4D semantic occupancy forecasting metrics of mIOU, IOU, and inference efficiency. Our approach achieves the best results in terms of both efficiency and performance.
@@
We investigate pretraining and fine-tuning a LiDAR world model across three diverse transfer tasks (Figure 1): varying-beam occupancy forecasting, indoor occupancy forecasting, and semantic occupancy forecasting. First, LiDAR hardware varies in beam count and scan patterns, often degrading model generalization zhang2023copilot4d ; ran2024towards . We consider this cross-sensor setting to assess robustness to hardware variations. Second, robots equipped with LiDAR sensors often operate in vastly different environments wilson2023argoverse ; zhou2021lidar ; han2024dr , from outdoor to indoor settings. While cross-domain adaptation has been extensively studied in perception tasks li2024towards ; michele2024train , it remains underexplored in the context of dynamic learning. Given that the depth range of LiDAR data in certain environments is scarce, it is promising to investigate whether dynamic knowledge learned in data-rich and large-depth-scope domains can be transferred to those with limited data availability and different absolute range. Lastly, for semantic world model tasks (e.g., semantic occupancy forecasting bian2024dynamiccity ; gu2024dome ; li2024uniscene ; zheng2024occworld ) which rely heavily on costly semantic labels wang2023openoccupancy ; tian2023occ3d . Our goal is to leverage large-scale unlabeled data to learn a universal 3D dynamics prior, enabling strong semantic forecasting via fine-tuning on minimal labeled data.
@@
Extensive experiments show that our LiDAR world model can significantly improve the convergence speed of downstream tasks: for all of the mentioned tasks, we observed relative performance gains at different amounts of fine-tuned data. Especially for the task of semantic occupancy forecasting, this scheme of learning environment dynamics before learning semantic patterns allows us to achieve better performance than OccWorld zheng2024occworld with only 5% of the labeled data. Furthermore, we note the importance of representation alignment in the fine-tuning tasks: Although our experiments demonstrate that the data compression structure we propose below is data-efficient, using either the pretrained data compressor directly or retraining the VAE from scratch with the fine-tuned data leads to underwhelming performance. This can be explained by subspace similarity: in both cases, the encoder learns a different feature space mapping from the pretrained one, which hurts the performance of flow models pre-trained on the original feature space. We observe two superior methods: fine-turning the VAE, and separately, applying a cosine-similarity-based alignment loss for tasks on which the original VAE can’t be fine-turned.
@@
We also show current architectural paradigms used in LiDAR world models peebles2023scalable ; esser2024scaling suffer from 2 issues: redundant model parameters and excessive training time. Regarding redundant model parameters: the latent representation tends to retain a large number of channels, which significantly increases dynamic learning model parameter counts. Regarding excessive training time: state-of-the-art models often require thousands of training epochs to converge. This inefficiency stems not only from model scale, but also from the inherently slow and compute-intensive nature of denoising diffusion paradigms—particularly those combining DDPM-based training with DDIM-style sampling.
@@
To address these issues, we propose a Swin Transformer-based VAE architecture for LiDAR data compression, achieving a compression ratio of 192x—over 6x higher than previous state-of-the-art methods—while always maintaining and sometimes exceeding reconstruction quality. We also propose a practical flow-matching-based generative model. Compared to previous latent diffusion-based schemes or transformer-based deterministic schemes, our approach requires only 4.38% and 28.91% of the FLOPs required by them respectively. In summary, our contributions are:
@@
  • To the best of our knowledge, this is the first study on building foundational LiDAR world models: world models of LiDAR videos that exhibit substantial transferability to downstream forecasting tasks. We show the efficacy of LiDAR world models to 3 diverse fine-tuning tasks: semantic occupancy forecasting, indoor occupancy forecasting, and high-beams occupancy forecasting, and confirm that it outperforms the baseline of training from scratch on the fine-tuning data, and that the relative performance gains is more pronounced with less fine-tuning data.
@@
  • With the decomposed dynamic pretrain + aligned semantic representation fine-tuning approach, we can significantly reduce reliance on human labeled samples for semantic occupancy forecasting: our method can exceed previous methods with only 5% of the semantic data.
@@
  • We design efficient architectures for data compression (VAE) and voxel-based LiDAR world modeling. The former achieves the SOTA reconstruction accuracy based on the 6x improvement in compression rate over prior work. Based on such a latent, our dynamic learning model achieves SOTA performance with only 4.47% to 50.23% of the FLOPs and 1.1 to 28.2 times higher FPS.
@@ 2 Related Work
Semantic 4D Occupancy Forecasting: The RGB video based world models tried to forecast physical consistent video from RGB input. However, these representations lack of geometric and explicit semantic annotation, the usage of the generated forecasts observations in control task still need extra component to recover depth and predict semantics. Semantic 4D occupancy forecasting aims to address this gap by predicting future semantic (LiDAR) occupancy maps based on past observations—either ground-truth annotations or model-generated results. OccWorld zheng2024occworld employed an auto-regressive transformer to jointly forecast future latent states and corresponding axis offsets. OccSora wang2024occsora further advanced the field by being the first to generate 25 seconds semantic videos conditioned on 512x compressed inputs. Later, DynamicCity bian2024dynamiccity improved upon this by decomposing the 3D representation into a more compact HexPlanecao2023hexplane structure, enabling faster inference. Building upon future trajectories/bev layout and extended datasets, DOME gu2024dome and uniScenes li2024uniscene continue to push the forecasting accuracy to new state-of-the-art levels. Importantly, these methods rely on labeled semantic ground-truth data for training, making them dependent on expensive human annotations and thus challenging to scale.
@@ 3.1 Data compression
Figure 2: Architecture of our VAE for LiDAR compression. The model enables high compression ratios—exceeding those of previous methods—alongside high-fidelity reconstructions.
@@
For compressing image-based RGB signals, SD3 esser2024scaling provides a strong baseline. However, LiDAR-specific data compression remains underexplored. Existing LiDAR world models typically adopt SD3’s encoder-decoder architecture with minimal modifications, retraining it directly on LiDAR BEV representations. Based on the Swin Transformer, we propose a new option and find it has the leading performance on all forms of LiDAR compression in Table 2. As noted in previous works li2024uniscene , discrete coding-based compression is not only prone to problems such as codebook collapse, but is also generally inefficient in compression. Here we have also used continuous coding to achieve a higher compression ratio as shown in Fig. 2.
@@
Data processing and encoding: Hereafter, bolding indicates a multidimensional array (e.g., a vector or a matrix). Given a raw LiDAR sensor scan $\bm{o}_{l}\in\mathbb{R}^{L\times 3}$, the voxelized map $\bm{o}_{v}$, dense occupancy map $\bm{o}_{d}$, and semantic occupancy map $\bm{o}_{s}$ can be generated through successive steps of voxelization, densification (using the entire LiDAR sequence to include points to each frame), and semantic annotation. Our structure is capable of processing any of these representations derived from the raw LiDAR input. Depending on the final reconstruction requirements, voxelization can be either a learning-based approach or a direct binarization process. Taking the compression of $\bm{o}_{d}\in\mathbb{R}^{H\times W\times D\times C}$(a $H,W,D$ 3D grid with C-dim features at each voxel) as an example, a Gaussian encoding distribution $q(\bm{z}_{d}|\bm{o}_{d})\!=\!\mathcal{N}(\bm{z}_{d};\bm{\mu}_{q},\bm{\sigma}_{q})$ is constructed from three main stages: embedding, feature learning and downsampling, and channel reduction. Specifically, after applying height and class embeddings, a 2D BEV feature map is obtained, which is subsequently processed by a standard 2D Swin Transformer encoder. Unlike the original Swin Transformer design, which utilizes patch merging for downsampling, we replace it with conventional convolutional layers, which will later be shown to outperform the former approach. Channel reduction is performed by a lightweight network neck, compressing the feature representation into a 16-dimensional latent space. Samples are drawn using the reparameterization trick $\bm{z}\!=\!\bm{\mu}_{q}+\bm{\sigma}_{q}\odot\bm{\epsilon}$, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$, and $\odot$ is the Hadamard product.
@@
Decoding and $\bm{o}_{l}$ representation recovery. The decoder $q(\hat{\bm{o}}_{d}\mid\bm{z}_{d})$ is designed to mirror the encoder by starting with the same symmetric 2D block structure. Different from the previous method that used 3D blocks in the decoder to increase temporal feature consistency, we found that it’s actually have a negative effect in our structure, for both reconstruction and final forecasting result, as shown in the ablation study. Depending on the source representation, the reconstructed $\hat{\bm{o}}_{d}$ can be used to render the points with a differentiable ray rendering module or get the occupancy map by calculating a similarity score with the class embedding.
@@ 3.2 Forecasting with Conditional flow matching
With the VAE proposed, we are able to mitigate the parameter redundancy: most of the parameters of the existing methods come from the high dimensionality of the latents. To further make the training of the model more efficient, we present a new structure based on flow matching, which we show leads to SOTA performance in both forecasting accuracy and calculation efficiency as shown in Table 1. For a semantic occupancy forecast task, given $\bm{o}_{s}^{t_{0}:t_{2}}$, our VAE will compress these continuous frames to $\bm{z}_{s}^{t_{0}:t_{2}}\in\mathbb{R}^{(t_{2}-t_{0})\times H\times W\times 16}$. Given the $t_{1}$ is the middle index of these frames, our objective is to obtain the future latent $\bm{z}_{s}^{t_{1}:t_{2}}$ from the historical latent $\bm{z}_{s}^{t_{0}:t_{1}}$ using a flow-matching model $G_{\theta}$.
Figure 3: Architecture of our conditional velocity field predictor at time $t$. Historical frame latents are extracted via a frame-wise VAE encoder, and noised future latents are formed by injecting noise at timestep $t$. Latents are concatenated along the time dimension and passed through DiT blocks.
As shown in Figure 4, during the training stage, given a time point $t\in[0,1]$ represents the progress along the probability path from initial distribution(Standard Gaussian) to target distribution, the noised future latent $x_{t}$, the historical trajectory $J^{t_{0}:t_{1}}\in\mathbb{R}^{(t_{1}-t_{0})\times 3}$, and historical observation $\bm{o}_{s}^{t_{0}:t_{1}}$, the velocity field $u_{t}^{\theta}$ is regressed. Specifically, the noising added follows a linear interpolation method as shown in the equation 1, where $\sigma$ used to balance the scale of noise and latents and $\epsilon\sim\mathcal{N}(0,1)$. The $\bm{z}_{s}^{t_{0}:t_{1}}$ will be used as a part of condition in training: we concatenate the historical latents and noised future latents along time dimension to get the input of ${G}_{\theta}$, denoted by $\mathbf{z}_{\text{joint}}\in\mathbb{R}^{(t_{2}-t_{0})\times H\times W\times 32}$.
@@
Based on the spatial-temporal DiT structure proposed in previous methods ma2024latte ; li2024uniscene ; gu2024dome , we observe that the convergence speed remains suboptimal. Specifically, given $\mathbf{z}_{\text{joint}}$, the initial 3D convolutional layer increases the channel dimension to $C^{\prime}$, resulting in $\mathbf{z}^{\prime}_{\text{joint}}\in\mathbb{R}^{(t_{2}-t_{0})\times H\times W\times C^{\prime}}$. The width and height dimensions are then flattened before being fed into the spatial DiT, where multi-head attention (MHA) is applied and normalized using AdaLN. However, applying the same strategy in the temporal DiT would only fuse pixels along the temporal axis, limiting the temporal receptive field to just a single latent pixel(in a block). Therefore, we observed that most of the previous method like uniScenes li2024uniscene and DOME gu2024dome use 14-18 stacked blocks to ensure the temporal consistency, which is also one of the reasons of redundant parameters. While this design remains feasible in Latte ma2024latte , which operates on 3D video patches, our setting involves individual latents containing only one frame of information, making the learning of temporal dependencies more challenging. Fortunately, this issue can be effectively mitigated by simply adding a 3D convolutional layer with a larger receptive field after the spatial DiT. Furthermore, organizing the network in a UNet-style architecture, as opposed to a single-stride DiT backbone, is shown through later experiments to further enhance the forecast performance.
The overall training objective was design following the Rectified Flow liu2022flow :

### v2

@@ Abstract
LiDAR-based world models offer more structured and geometry-aware representations than their image-based counterparts. However, existing LiDAR world models are narrowly trained; each model excels only in the domain for which it was built. This raises a critical question: can we develop LiDAR world models that exhibit strong transferability across multiple domains? To answer this, we conduct the first systematic domain transfer study across three demanding scenarios: (i) outdoor to indoor generalization, (ii) sparse- to dense-beam adaptation, and (iii) non-semantic to semantic transfer. Given different amounts of fine-tuning data, our experiments show that a single pretrained model can achieve up to 11% absolute improvement (83% relative) over training from scratch and outperforms training from scratch in 30/36 of our comparisons. This transferability significantly reduces the reliance on manually annotated data for semantic occupancy forecasting: our method exceeds previous baselines with only 5% of the labeled training data of prior work. We also observed inefficiencies of current generative-model-based LiDAR world models, mainly through their under-compression of LiDAR data and inefficient training objectives. To address these issues, we propose a latent conditional flow matching (CFM)-based framework that achieves state-of-the-art reconstruction accuracy using only half the training data and a compression ratio 6 times higher than that of prior methods. Our model also achieves SOTA performance on semantic occupancy forecasting while being 1.98x-23x more computationally efficient (a 1.1x-3.9x FPS speedup) than previous methods. Our project page contains additional visualizations and released code.
@@ 1 Introduction
World models enable agents to implicitly learn the dynamics of the environment by predicting future sensory observations, typically through generative models operating in a latent space rombach2022high ; ma2024latte . Starting from classic motion prediction tasks like optical flow long2024learning ; long2022detail , recent advances have extended motion prediction to RGB-video-based forecasting, i.e., RGB world models, which have demonstrated impressive performance in applications such as autonomous driving russell2025gaia ; hu2023gaia ; wang2024driving , robotic navigation bar2024navigation , and other embodied tasks black2024pi_0 ; zhang2025mem2ego ; liu2025omnieva ; zhang2024plan ; kim2024openvlaopensourcevisionlanguageactionmodel ; agarwal2025cosmos ; alhaija2025cosmos . These models can generate video sequences conditioned on historical frames and sometimes natural language input. While an image-based world model may suffice for repetitive tasks (e.g., learning robotic arm movements), its utility is limited in complex scenarios that demand geometrically structured information, such as autonomous driving. In these scenarios, the lack of explicit semantic and geometric representations restricts practical application, as it necessitates additional steps to extract spatial information.
@@
In contrast to RGB images, LiDAR provides rich geometric structure and implicitly offers a sparse representation for semantic cues about the environment. Unlike dense, pixel-based observations, a 3D object in a LiDAR point cloud is represented as a cluster of 3D points, encoding essential spatial and semantic information in a compact form. Despite these advantages, foundational world models of LiDAR data remain relatively underexplored: prior LiDAR world models were built for specific domains zhang2023copilot4d ; khurana2023point ; zhang2024bevworld . In comparison, recent advances in “foundation” models, e.g. pretraining with RGB image forecasting alhaija2025cosmos and trajectory forecasting zhou2024smartpretrain have demonstrated the effectiveness of pretraining in improving performance on downstream tasks. This motivates our central question: Can we develop a foundational LiDAR world model for ground vehicles that yields downstream performance gains on diverse forecasting tasks, particularly those suffering from data limitations, after fine-tuning? This goal is partly motivated by the fact that many of the causal factors that govern the motion of objects are shared across different domains and environments—many aspects of the dynamics of the world should, in principle, be transferable and can be observed in unlabeled LiDAR.
@@
Figure 1: Left: The overall pipeline of our method: we used the most readily publicly available LiDAR dataset from autonomous driving scenarios to train the proposed LiDAR world model. This well-trained world model is able to generalize well on the listed downstream tasks after fine-tuning, although scene and signal properties are quite different. Right: Comparison of our proposed world model with previous methods on nuScenes 4D semantic occupancy forecasting metrics of mIoU, IoU, and inference efficiency. Our approach achieves the best results in terms of both efficiency and performance.
@@
We investigate pretraining and fine-tuning a LiDAR world model across three diverse transfer tasks (Fig  1): varying-beam occupancy forecasting, indoor occupancy forecasting, and semantic occupancy forecasting. First, LiDAR hardware varies in beam count and scan patterns, often degrading model generalization zhang2023copilot4d ; ran2024towards . We consider this cross-sensor setting to assess robustness to hardware variations. Second, robots equipped with LiDAR sensors often operate in vastly different environments wilson2023argoverse ; zhou2021lidar ; han2024dr , from outdoor to indoor settings. While cross-domain adaptation has been extensively studied in perception tasks li2024towards ; michele2024train ; long2025shape , it remains underexplored in the context of dynamic learning. Given that LiDAR data can be scarce in certain environments (e.g., indoor), it is promising to investigate whether dynamic knowledge learned in data-rich, long-range domains can be transferred to those with limited data and different operational ranges. Lastly, tasks like semantic occupancy forecasting bian2024dynamiccity ; gu2024dome ; li2024uniscene ; zheng2024occworld ) rely heavily on costly semantic labels wang2023openoccupancy ; tian2023occ3d . Our goal is to leverage large-scale unlabeled data to learn a universal 3D dynamics prior for ground vehicles, enabling strong semantic forecasting via fine-tuning on minimal labeled data.
@@
Extensive experiments demonstrate that our LiDAR world model significantly improves the convergence speed of downstream tasks. For all mentioned tasks, we observed relative performance gains with varying amounts of fine-tuning data. In particular, for semantic occupancy forecasting, this scheme of learning dynamics prior to semantics patterns allows us to achieve superior performance to OccWorld zheng2024occworld with only 5% of its required labeled data. Furthermore, our results highlight the importance of representation alignment during fine-tuning. Although the data compression structure we propose later is data-efficient, we found that either using the pretrained data compressor directly or retraining the VAE from scratch with the fine-tuning data leads to suboptimal performance. This phenomenon is attributable to a feature space mismatch: in both scenarios, the encoder learns a feature space mapping that diverges from the pretrained one, which in turn degrades the performance of the flow model pretrained on the original feature space. We identify two superior strategies: fine-tuning the VAE, and for tasks where the VAE cannot be fine-tuned directly, applying a cosine-similarity-based alignment loss.
@@
We also find that the current architectural paradigms used in LiDAR world models peebles2023scalable ; esser2024scaling suffer from 2 issues: redundant model parameters and excessive training time. Regarding the former, the latent representation tends to retain a large number of channels, which significantly increases the parameter count of the dynamics model. Regarding the latter, state-of-the-art models often require thousands of epochs to converge. This inefficiency stems not only from the model scale but also from the inherently slow and compute-intensive nature of denoising diffusion paradigms—particularly those combining DDPM-based training with DDIM-style sampling. These challenges have significantly hindered our exploration of LiDAR world model transferability.
@@
To address these issues, we first propose a Swin Transformer-based VAE architecture for LiDAR data compression. This architecture achieves a compression ratio of 192x—over 6x higher than previous state-of-the-art methods—while matching or exceeding their reconstruction quality. We also propose an efficient flow-matching-based generative model. Compared to previous latent diffusion-based or transformer-based deterministic schemes, our approach requires only 4.38% and 28.91% of the FLOPs that they require, respectively. This efficiency significantly accelerates our analysis of transferability in downstream tasks. In summary, our contributions are:
@@
  • The first study on building transferable LiDAR world models: world models of LiDAR videos that exhibit substantial transferability to downstream forecasting tasks. We show the efficacy of LiDAR world models to 3 diverse fine-tuning tasks: semantic occupancy forecasting, indoor occupancy forecasting, and high-beams occupancy forecasting, and confirm that it outperforms the baseline of training from scratch on the fine-tuning data, and that the relative performance gain is more pronounced with less fine-tuning data.
@@
  • An approach to substantially reduce reliance on human labels for semantic occupancy forecasting: our method exceeds the performance of previous methods with only 5% of the human labels.
@@
  • Efficient VAE-baesd architectures for data compression and voxel-based LiDAR world modeling. The former achieves the SOTA reconstruction accuracy at a 6x improvement in compression rate over prior work. Using these encodings, our world model achieves SOTA performance with only 4.47% to 50.23% of the FLOPs and 1.1 to 3.9 times higher FPS.
@@ 2 Related Work
Semantic 4D Occupancy Forecasting: The RGB video-based world models tried to forecast physical consistent video from RGB input. However, these representations lack of geometric and explicit semantic annotation, the usage of the generated forecasts observations in control task still need extra component to recover depth and predict semantics. Semantic 4D occupancy forecasting aims to address this gap by predicting future semantic (LiDAR) occupancy maps based on past observations—either ground-truth annotations or model-generated results. OccWorld zheng2024occworld employed an auto-regressive transformer to jointly forecast future latent states and corresponding axis offsets. OccSora wang2024occsora further advanced the field by being the first to generate 25 seconds semantic videos conditioned on 512x compressed inputs. Later, DynamicCity bian2024dynamiccity improved upon this by decomposing the 3D representation into a more compact HexPlanecao2023hexplane structure, enabling faster inference. Building upon future trajectories/bev layout and extended datasets, DOME gu2024dome and uniScenes li2024uniscene continue to push the forecasting accuracy to new state-of-the-art levels. Importantly, these methods rely on labeled semantic ground-truth data for training, making them dependent on expensive human annotations and thus challenging to scale.
@@ 3.1 Data compression
Figure 2: Architecture of our VAE for LiDAR compression. The model enables high compression ratios—exceeding those of previous methods—alongside high-fidelity reconstructions. Here we use $\bm{O}_{v}$ to represent the raw voxelized LiDAR points (Occupancy) and $\bm{O}_{d}$, $\bm{O}_{s}$ stand for densified or semantic labeled occupancy respectively.
@@
While SD3 esser2024scaling provides a strong baseline for compressing image-based RGB signals, LiDAR-specific data compression remains underexplored. Existing LiDAR world models typically adopt the SD3 encoder-decoder architecture with minimal modifications, retraining it directly on LiDAR BEV representations. As an alternative, we propose a Swin Transformer-based architecture and demonstrate that it achieves state-of-the-art performance across all forms of LiDAR compression, as detailed in Table 2. As noted in previous works li2024uniscene , discrete coding-based compression is not only prone to problems such as codebook collapse but is also generally inefficient in compression. Here we have also used continuous coding to achieve a higher compression ratio as shown in Figure 2.
@@
Data processing and encoding: In this work, bold notation indicates a multidimensional array (e.g., a vector or a matrix). Given a raw LiDAR sensor scan $\bm{o}_{l}\in\mathbb{R}^{L\times 3}$, the voxelized map $\bm{o}_{v}$, dense occupancy map $\bm{o}_{d}$, and semantic occupancy map $\bm{o}_{s}$ can be generated through successive steps of voxelization, densification (i.e., accumulating points from the entire sequence into each frame), and semantic annotation. Our structure is capable of processing any of these representations derived from the raw LiDAR input. Depending on the final reconstruction requirements, voxelization can be either a learning-based approach or a direct binarization process. Taking the compression of $\bm{o}_{d}\in\mathbb{R}^{H\times W\times D\times C}$ (a $H,W,D$ 3D grid with C-dim features at each voxel) as an example, a Gaussian encoding distribution $q(\bm{z}_{d}|\bm{o}_{d})\!=\!\mathcal{N}(\bm{z}_{d};\bm{\mu}_{q},\bm{\sigma}_{q})$ is constructed from three main stages: embedding, feature learning and downsampling, and channel reduction. Specifically, after applying height and class embeddings, a 2D BEV feature map is obtained, which is subsequently processed by a standard 2D Swin Transformer encoder. Unlike the original Swin Transformer design, which utilizes patch merging for downsampling, we replace this operation with conventional convolutional layers—a modification we empirically find to outperform the original approach. Channel reduction is performed by a lightweight network neck, compressing the feature representation into a 16-dimensional latent space. Samples are drawn using the reparameterization trick $\bm{z}\!=\!\bm{\mu}_{q}+\bm{\sigma}_{q}\odot\bm{\epsilon}$, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$, and $\odot$ denotes the Hadamard product.
@@
Decoding and $\bm{o}_{l}$ representation recovery. The decoder $q(\hat{\bm{o}}_{d}\mid\bm{z}_{d})$ is designed to mirror the encoder by starting with the same symmetric 2D block structure. Different from previous methods that used 3D blocks in the decoder to increase temporal feature consistency, we found doing so to have a negative effect in our structure, for both reconstruction and final forecasting result, as shown in the ablation study. Depending on the source representation, the reconstructed $\hat{\bm{o}}_{d}$ can be used to render the points with a differentiable ray rendering module or get the occupancy map by calculating a similarity score with the class embedding.
@@ 3.2 Forecasting with Conditional flow matching
With the proposed VAE, we are able to mitigate the parameter redundancy: most of the parameters of the existing methods come from the high dimensionality of the latents. To further make model training more efficient, we present a new structure based on flow matching liu2022flow , which we show leads to SOTA performance in both forecasting accuracy and computational efficiency, as shown in Table 1. For a semantic occupancy forecast task, given $\bm{o}_{s}^{t_{0}:t_{2}}$, our VAE encodes continuous frames to $\bm{z}_{s}^{t_{0}:t_{2}}\in\mathbb{R}^{(t_{2}-t_{0})\times H\times W\times 16}$. As $t_{1}$ is the middle index of these frames, our objective is to obtain the future latent $\bm{z}_{s}^{t_{1}:t_{2}}$ from the historical latent $\bm{z}_{s}^{t_{0}:t_{1}}$ using a flow-matching model $G_{\theta}$.
Figure 3: Architecture at the training stage of our proposed conditional velocity field predictor for time $t$. Historical frame latents are extracted via a frame‐wise VAE encoder, and noised future (target) latents are formed by injecting noise at timestep $t$. Latents are concatenated along the time dimension and passed through DiT blocks.
As illustrated in Figure 3, the training objective is to regress a velocity field $u_{t}^{\theta}$. This regression is conditioned on several inputs: a time step $t\in[0,1]$ representing the progress along a probability path from an initial distribution (standard Gaussian) to the target distribution; the interpolated future latent state $x_{t}$; the historical trajectory $J^{t_{0}:t_{1}}\in\mathbb{R}^{(t_{1}-t_{0})\times 3}$; and historical observations $\bm{o}_{s}^{t_{0}:t_{1}}$.
Specifically, the noising added follows a linear interpolation method as shown in equation 1, where $\sigma$ used to balance the scale of noise and latents and $\epsilon\sim\mathcal{N}(0,1)$. The $\bm{z}_{s}^{t_{0}:t_{1}}$ will be used as a part of condition in training: we concatenate the historical latents and noised future latents along time dimension to get the input of ${G}_{\theta}$, denoted by $\mathbf{z}_{\text{joint}}\in\mathbb{R}^{(t_{2}-t_{0})\times H\times W\times 32}$.
@@
Based on the spatial-temporal DiT structure proposed in previous methods ma2024latte ; li2024uniscene ; gu2024dome , we observe that the convergence speed remains suboptimal. Specifically, given $\mathbf{z}_{\text{joint}}$, the initial 3D convolutional layer increases the channel dimension to $C^{\prime}$, resulting in $\mathbf{z}^{\prime}_{\text{joint}}\in\mathbb{R}^{(t_{2}-t_{0})\times H\times W\times C^{\prime}}$. The width and height dimensions are then flattened before being fed into the spatial DiT, where multi-head attention (MHA) is applied and normalized using AdaLN. However, applying the same strategy in the temporal DiT would only fuse pixels along the temporal axis, limiting the temporal receptive field to just a single latent pixel(in a block). To compensate, we observed that most of the previous methods like uniScenes li2024uniscene and DOME gu2024dome use 14-18 stacked blocks to ensure the temporal consistency, which is also one of the reasons of redundant parameters. While this design remains feasible in Latte ma2024latte , which operates on 3D video patches, our per-frame latent representation, making the learning of temporal dependencies more challenging. We found that this issue can be effectively mitigated by simply inserting a 3D convolutional layer with a larger receptive field after the spatial DiT. Furthermore, organizing the network in a UNet-style architecture, as opposed to a single-stride DiT backbone, is shown through later experiments to further enhance the forecast performance and reduce computational load. The training objective was designed following the Rectified Flow liu2022flow :